Calculate the molarity of a solution prepared by dissolving 250.0 g of Ca(NO3)2 in enough water to make 3.50 L of solution.

 a

5.34 M

 b

0.45 M

 c

1.52 M

 d

71.4 M


1
Expert's answer
2022-04-07T04:07:39-0400

M(Ca(NO3)2=40+2(14+16x3))=40+144=164 g/mol

n(Ca(NO3)2)=250/164=1.524mol

c(Ca(NO3)2)=1.524/(3.5)=0.45 mol/l